OpenStudy (inowalst):

Four times the first of three consecutive even integers is six more than the product of two and the third integer. Find the integers.

OpenStudy (badhi):

I think there is something missing here you need to have and operator for the first three consecutive even integers. Is it four times the "sum of" three consecutive even integers or is it four times the "multiplication" of the consecutive even integers?

OpenStudy (texaschic101):

3 consecutive EVEN integers... x, x + 2, x + 4 4(x) = 2(x + 4) + 6 4x = 2x + 8 + 6 4x - 2x = 8 + 6 2x = 14 x = 7 must be doing something wrong because I am getting an odd number

OpenStudy (texaschic101):

@phi ...help set this up please

OpenStudy (inowalst):

Yes, but the integers are x, x+2, x+4

OpenStudy (texaschic101):

but x has to be an even number...not an odd one

OpenStudy (inowalst):

Hm.

OpenStudy (texaschic101):

@Loser66 ...help

OpenStudy (inowalst):

@paki

OpenStudy (texaschic101):

we WILL figure this out...lol

OpenStudy (inowalst):

lol Okay

OpenStudy (texaschic101):

phi...what did I screw up on ?

OpenStudy (inowalst):

x=7 x+2=7+2=9 x+4=7+4=11 ???

OpenStudy (texaschic101):

but the integers need to be even...not odd

OpenStudy (phi):

I think the problem is "ill-posed"

OpenStudy (loser66):

You need form the even integer as 2n, 2n+2, 2n+4 , not x

OpenStudy (loser66):

hence "four times the first of them" is 4*2n =8n product of the two and the third is (2n+2)(2n+4) blahblah....

OpenStudy (texaschic101):

wait....is it the produce of 2 and the 3rd integer....or is it the product of the 2nd and 3rd integer ?

OpenStudy (inowalst):

I think its of 2 and the 3rd integer.

OpenStudy (acxbox22):

yeah it cant be 2nd integer...i tried it already

OpenStudy (texaschic101):

4(2n) = 2(2n + 4) + 6 8n = 4n + 8 + 6 8n - 4n = 14 4n = 14 nope....I am not understanding this

OpenStudy (texaschic101):

unless the integers are fractions ?? ....sorry....I do not know this one

OpenStudy (texaschic101):

nvm..integers cannot be fractions

OpenStudy (loser66):

I think: "No solution" is a solution.

OpenStudy (texaschic101):

oh.....thats tricky

OpenStudy (acxbox22):

are you sure it isnt consecutive ODD integers?

OpenStudy (texaschic101):

now that would work...lol

OpenStudy (loser66):

Suppose it is 2, 4,6 how can 4*2 = (4)(6) + 6??

OpenStudy (texaschic101):

true ^^

OpenStudy (acxbox22):

or 12,14,16 4 x 12 = (14 x 16) +6? or 4 x 12= (2x16)+6

OpenStudy (theraggedydoctor):

Okay.